cobalt crystallizes in a HCP unit cell. it's atomic radius is 0.1253 nm. it's density is 8900 kg/m^3. using this information, estimate avogadro's number.


1
Expert's answer
2021-04-28T06:18:41-0400

Density= 8900Kg/m³= 8.9g/cm


Radius= 0.1253nm= 1.253x10-8


One mole of Co= 59g


Volume occupied by one mole of Co= mass/density= 59/8.9 = 6.63cm3


Volume of one atom of Co= 4/3Ï€r3


= 4/3 x 3.142x (1.253x10-8)³


= 8.24x10-24cm3


number of atoms present in one mole of Cobalt= volume occupied by one mole/volume occupied by one atom


= 6.63/8.24x1023


= 8.04x1023 atoms


We can see that this value is close to the Avogadro's number which is 6.02x1023 atoms.